Solve the following system using any method you like: 3x – 2y = 1 –2x + 2y = 1
Answer #1

*> Here, the system is 371-2y = -22 +2y = 1 2 0+® => 2 = 1+1 = 2 3 Therefore, using using @ we have 2y = 1+2x = 1+ 2(2) = 5 [Here is the required solution.Hope this helps you.Please upvote my work if you get benefited.Your feedback is very much precious.Thank You.